# Training in reverse: How iteration order influences convergence and stability in deep learning

### v1

@@ 2.1 A fundamental example of a contraction
We are now describing how the SGD updates are uniform contractions when the batch losses are strictly convex near a minimum. First, consider a strictly convex smooth function $L$. Recall that $L$ is strictly convex if 1) there is a constant $m>0$ such that
$\langle\nabla L(\theta_{1})-\nabla L(\theta_{2}),\theta_{1}-\theta_{2}\rangle\geq m\left\|\theta_{1}-\theta_{2}\right\|^{2},$
where $\langle a,\,b\rangle=a^{T}b$ is the inner product between two vectors, and 2) there is a constant $M>m$ such that
$\left\|\nabla L(\theta_{1})-\nabla L(\theta_{2})\right\|\leq M\left\|\theta_{1}-\theta_{2}\right\|.$
The following lemma (whose proof is in Appendix F) shows that for strictly convex functions we can always find a learning rate small enough so that the corresponding gradient descent update is a contraction.
@@
Let $L(\theta)$ be a strictly convex function and define the map $T(\theta)=\theta-h\nabla L(\theta)$. Then
@@
Now if we suppose that batch losses $L_{i}$’s are strictly convex (possibly near global minima), then Lemma 2.5 tells us that the SGD maps $T_{i}(\theta)=\theta-h\nabla L_{i}(\theta)$ are uniform contractions. Thus we can apply Theorem 2.2 and see that the sequence of backward iterates, $\theta_{n}=T_{1}T_{2}\cdots T_{n}(\theta)$, converges to a point.
2.2 Relation between forward and backward iterations
As seen in Example 2.4, an actual realization of a forward trajectory converges to a distribution which is uniformly distributed between the two outcomes. On the other hand, any realization of the backward trajectory always converges to a point sampled from this distribution. Let’s formalize this for the general case.
@@ 2.2 Relation between forward and backward iterations
Therefore, the forward sequence $(T_{n}\cdots T_{2}T_{1}):\Omega\rightarrow\Omega$ induces a push-forward $(T_{n}\cdots T_{2}T_{1})_{*}$ that maps an initial distribution $\mu_{0}$ to the probability measure $\mu_{n}=(T_{n}\cdots T_{2}T_{1})_{*}\mu_{0}$ modelling the probability distribution of the $n^{th}$ forward iterate $\theta_{n}=T_{n}\cdots T_{2}T_{1}(\theta_{0})$ when the initial point $\theta_{0}$ is randomly sampled from $\mu_{0}$. Observe that the forward iterates $\theta_{n}$ form a sequence of random variables, which is a Markov chain (since for the forward iterates we have that $\theta_{n}=T_{n}(\theta_{n-1})$). Many previous works (see related work in Section 1.1) have shown under different assumptions that this sequence of forward iterates do not converge to a point when using fixed learning rates, but rather their probability distributions $\mu_{n}$ converge to a limiting probability distribution, called the stationary distribution of the Markov chain. In other words, starting at initial point $\theta_{0}$ the distribution $\mu_{n}=(T_{n})_{*}\cdots(T_{1})_{*}\delta_{\theta_{0}}$ of the forward iterate $\theta_{n}$ converges to a stationary distribution $\mu_{\theta^{*}}$. Recall that $\delta_{\theta_{0}}$ is the Dirac measure concentrated at $\theta_{0}$ (i.e., $\delta_{\theta_{0}}(A)$ is $1$ if $A$ contains $\theta_{0}$ and 0 otherwise).
On the contrary, as we have seen in Theorem 2.2, the backward iterates converge to actual points in regions where the maps $T_{i}$’s become contractions (which we will call contractive regions). The following theorem (whose proof is in Appendix E) shows the relation between the point-wise convergence of the backward iterates and the distributional convergence of the forward iterates:
@@
Consider a sequence $\{T_{i}\}$, $i=1,2,\dots$ of independent and identically distributed random operators. Suppose that for $\theta_{0}\in\Omega$ the backward iterates converge to a random point (randomness is due to the sampling of the random operators $T_{i}$’s):
@@ 3.1 Example: Quadratic loss
Consider the quadratic loss function $L(\theta)=\theta^{2}/2$ with $\theta\in\mathbb{R}$. Its gradient is $\theta$. We want to find the minimum of $L$ (which of course is 00 at $\theta=0$) by a stochastic gradient descent. To model the batch noise, we add a i.i.d random error $\epsilon_{i}$ to the gradient at step $i$, yielding the iteration procedure:
@@
$T_{n}\cdots T_{1}(\theta)=(1-h)^{n}\theta+h\epsilon_{1}+\cdots+h(1-h)^{n-1}\epsilon_{n}.$
@@ 3.2 Example: Learning dynamics close to minima
$L(\theta)=\frac{1}{N}\sum_{i=1}^{N}L_{i}(\theta).$
@@
since at a minimum $g(\theta^{*})=0$. To simplify the notation, we will denote by $g_{i}(\theta)$ the gradient of the batch loss $L_{i}(\theta)$, which is now a random vector (because the batch is random). We can model this stochastic gradient as follows: $g_{i}(x)=g(x)-\epsilon_{i}$, where we assume that the differences $\epsilon_{i}=g(x)-g_{i}(x)$ are i.i.d random vectors with zero mean. Using gradient approximation (3) we have that a single gradient update is given in a form very similar to that of the previous example:
$T_{i}(\theta)=\theta-h\Big{(}g(\theta)-\epsilon_{i}\Big{)}\simeq\theta^{*}+(1-hH)(\theta^{*}-\theta)+h\epsilon_{i}$
@@
and since $\|1-hH\|<1$ for $h$ strictly smaller than $2/\lambda_{max}$ (as in Example 2.1), Theorem 2.2 tells us that the backward trajectory converges. Now let us have a look at the form of the forward and backward iterates in this particular case to give us an intuition about why this is true. In a similar way than in the previous example, the $n^{th}$ forward SGD iterate is given by
$T_{n}\cdots T_{1}(\theta)$ $=$ $\theta^{*}+(1-hH)^{n}(\theta-\theta^{*})$
$+h\epsilon_{n}+h(1-hH)\epsilon_{n-1}+\cdots$
$\cdots+h(1-hH)^{n-1}\epsilon_{1}.$
While the term $(1-hH)^{n}(\theta-\theta^{*})$ converges to zero (when $h<2/\lambda_{max}$), we see that at each step $n$ a fresh random vector $h\epsilon_{n}$ is added to the iterate making the whole sequence converge to a distribution rather than to a point. On the other hand the backward iterate is obtained by reversing the indices:
$T_{1}\cdots T_{n}(\theta)$ $=$ $\theta^{*}+(1-hH)^{n}(\theta-\theta^{*})$
$+h\epsilon_{1}+h(1-hH)\epsilon_{2}+\cdots$
$\cdots+h(1-hH)^{n-1}\epsilon_{n}.$
This expression converges as $n\rightarrow\infty$ since the terms are decreasing at an exponential rate as $n$ grows. Thus the iterates converge to a point now rather than a distribution. Looking at the actual form of the backward iterates, it is clear that they are distributed by the probability measure given by the forward iterates as prescribed by Theorem 2.6.
@@ 4 Experiments
Note on plotting multiple seeds: We are interested in the variability per realization of the backward and forward trajectories. The backward trajectories are more stable individually along their own path, but these paths can be very different from seed to seed (because of the convergence toward different points). Therefore the phenomenon is much clearer when the seeds are plotted individually rather than averaged, which creates artificially more variability in the backward trajectories as there really is on each individual realization. This is why we are reporting the various seed plots individually and not as a single averaged plot with error bars. Note that the increased stability and point convergence is visible in each of the seeds, which are added to the Appendix B.
@@
We trained a ResNet-18 with stochastic gradient descent and no regularization on the CIFAR-10 dataset [44]. We used a learning rate of 0.025 and a batch-size of 8. In Figure 2, we recorded the learning curves at each gradient update for both the forward and backward iterations for 2000 steps. The additional seeds are in Section B.1. We observe that the training loss for the backward trajectory is more stable, converges faster, and has less variability than the forward trajectory, and similarly for all the other learning curves. In Appendix A, we repeat the experiment with different datasets (synthetic dataset, FashionMNIST, and CIFAR-100) as well as different architectures (ResNet-50, VGG, and MLP). Each time we observe the same phenomenon for all 5 seeds.
@@
We trained a MLP with 5 layers of 500 neurons each with stochastic gradient descent with no regularization to learn FashionMNIST [45]. The batch-size was set to 8 while the learning rate was 0.001. In Figure 3, we report the test accuracy at every single step for 2000 steps. At step 1000, we reset the point from which we perform backward SGD to that of the parameter at this step. As a result we see that the backward trajectory from that reset point seems to converge again, but to a different point (with different test accuracy), while forward SGD seems to oscillate between these two points. This is in line with the theoretical prediction that backward trajectories SGD converge toward points distributed according to the distribution induced by forward trajectories (see Theorem 2.6). The additional seeds are in Section B.2.
Figure 2: Backward SGD exhibits decreased variance and increased stability compared to forward SGD for a ResNet-18 model trained on CIFAR-10. The additional seeds are in Section B.1.
Figure 3: Backward SGD converges toward a different minima after resetting the initialization point at step 1000 (“intermittent backward”) while forward SGD oscillates between them for MLP trained on FashionMNIST. Top: On the first seed, backward changes from a higher test-performance trajectory to a lower test-performance trajectory at the reset step 1000. Bottom: On the second seed, backward changes this time from a trajectory converging to a lower test-performance point to a trajectory converging to a higher test-performance point. The other seeds can be found in Appendix B.2 including all the learning curves.
@@ 5 Conclusion and future work
In this paper, we analyzed the impact of two different iteration orders on the convergence and stability of stochastic learning trajectories. Although the very idea of backward trajectories may seem strange and counterintuitive at first, we show that this approach of backward optimization has clear advantages. Namely, we show theoretically that the backward trajectories converge toward actual points while the forward trajectories converge toward probability distributions in regions where the optimization maps become contractions, leading to improved stability and convergence. We observed this phenomenon experimentally on a number of datasets (synthetic, FashionMNIST, CIFAR-10, and CIFAR-100) and neural network architectures (MLP, VGG-19, ResNet-18, and ResNet-50), making it relevant to deep learning. This phenomenon points toward the fact that the ordering in which the data examples are consumed to produce each parameter iterate impacts the properties of both the learning trajectories and their convergence points. In particular the last examples used to produce a parameter iterate seem to have great importance. Indeed, the backward order is the only order in which the sequence of the last examples used to produce a parameter iterate remains consistent at each training step. As we saw, this yields convergence toward a point rather than oscillations between solutions of varied performances.
@@
One possible strategy to mitigate the intensive computational requirements of the full backward order is to apply it only partially on a fixed window of iterations from a point that is reset periodically. This way the extra computation time is capped by the window size. We see that this approach leads to convergence and increased stability in Figure 3; however the backward trajectory may change and converge to a different point at each reset. We can imagine strategies to initiate this reset only when the performance of the new backward trajectory increases, leading to stable convergence to higher test performance points. Another possibility is to start the backward order only later in training to force convergence toward a single solution after the forward trajectory has done enough exploration of the space as demonstrated in Figure 10 in Appendix A. We imagine that this approach may possibly lead toward more reliable early-stopping criteria. Similarly, application of the backward order at the very beginning of training may give an accurate idea of whether a specific hyper-parameter setting is promising or needs to be aborted by creating stabler trajectories whose learning curves are easier to interpret early on.
An orthogonal approach to deal with the computational challenges of backward optimization is to approximate the backward trajectory. In Appendix C, we show how to compute a backward iterate from a cheaper forward iterate to which we add a correction that can be computed completely independently. We compute these corrections up to order 2 in Theorem C.2 Appendix C.2, namely:
@@
where $[\nabla L_{i},\nabla L_{j}](\theta)=H_{i}(\theta)\nabla L_{j}(\theta)-H_{j}(\theta)\nabla L_{i}(\theta)$ is the Lie bracket between the vector fields $\nabla L_{i}$ and $\nabla L_{j}$. Even though the approximation we provide is not immediately useful because the higher order terms seem to matter, it probably will be useful if higher order corrections are included. As a proof of the usefulness of these type of corrections added to the forward iterates to emulate alternate orderings, we show in Appendix D that the corrections we computed at second order are already enough to approximate an average between all the possible iteration orders on a batch window. Adding these corrections to the forward iterates produces a beneficial regularizer that mimics small batch training regularization. In fact, a training step can be seen as a parameter average of mixture of models where each model is trained not from a different seed but from a different batch order. We hope that this paper will bring awareness and foster more research toward understanding and exploiting the role of iteration order in the production of models with more consistent properties along more stable learning trajectories.
Acknowledgements
We would like to thank Michael Wunder, Peter Bartlett, George Dahl, Atish Agarwala, and Scott Yak for helpful discussion and feedback.

### v2

@@ 2.1 A fundamental example of a contraction
We now describe how the SGD updates are uniform contractions when the batch losses are strongly convex near a minimum. First, consider a strongly convex smooth function $L:\mathbb{R}^{d}\to\mathbb{R}$. Recall that $L$ is strongly convex if there is a constant $m>0$ such that for all $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$
$\langle\nabla L(\theta_{1})-\nabla L(\theta_{2}),\theta_{1}-\theta_{2}\rangle\geq m\left\|\theta_{1}-\theta_{2}\right\|^{2},$ (3)
where $\langle a,\,b\rangle=a^{T}b$ is the inner product between two vectors. Furthermore, we assume that there is a constant $M>m$ such that
$\left\|\nabla L(\theta_{1})-\nabla L(\theta_{2})\right\|\leq M\left\|\theta_{1}-\theta_{2}\right\|.$ (4)
This is sometimes denoted as $L\in\mathcal{S}_{m,M}^{1,1}(\mathbb{R}^{d})$, see Nesterov (2013). The following lemma (whose proof is in Appendix˜F) shows that for strongly convex functions we can always find a learning rate small enough so that the corresponding gradient descent update is a contraction.
@@
Let $L(\theta)$ satisfy ˜3 and 4 and define the map $T(\theta)=\theta-h\nabla L(\theta)$. Then
@@
Now if we suppose that batch losses $L_{i}$’s are strongly convex (possibly near global minima), then Lemma˜2.5 tells us that the SGD maps $T_{i}(\theta)=\theta-h\nabla L_{i}(\theta)$ are uniform contractions. Thus, we can apply Theorem˜2.2 and see that the sequence of backward iterates, $\theta_{n}=T_{1}T_{2}\cdots T_{n}(\theta)$, converges to a point.
2.2 Relation between forward, backward, and generalization
In this section we address the question of the generalization power of backward iterates, which we argue is the same as that of the forward iterates. The relationship with generalization is implicit through Theorem˜2.6 below which states that backward convergence points are distributed according to the forward stationary distribution. It follows that the best test-performance forward-solutions can be reached by the backward trajectories with the same probability. This a good thing as it shows that the backward trajectories are more stable but without test-performance reduction on average, which is often not the case as for instance the increased stability obtained by increasing the learning rate or decreasing the batch-size (see Dherin et al. (2022)). In particular, the bias of backward SGD is statistically equivalent to that of standard SGD. Consequently, the convergence point of the backward iterates cannot outperform the best solutions attainable by the forward iterates.
The goal of this work aims to demonstrate that the composition order creates two different processes, which we could described intuitively as follows:
  • The forward iterate process converges toward a distribution, “forgetting” the initial batches and being most influenced by the recent ones. This makes this process robust to a bad start but leaves it perpetually noisy.
  • The backward iterate process converges toward a point, “remembering” the initial batches but “forgetting” the recent stochastic noise.
Theorem˜2.6 formalizes this intuition and essentially generalizes the situation already seen in Example˜2.4, an actual realization of a forward trajectory converges to a distribution which is uniformly distributed between the two outcomes. On the other hand, any realization of the backward trajectory always converges to a point sampled from this distribution. Let’s formalize this for the general case.
@@ 2.2 Relation between forward, backward, and generalization
Therefore, the forward sequence $(T_{n}\cdots T_{2}T_{1}):\Omega\rightarrow\Omega$ induces a push-forward $(T_{n}\cdots T_{2}T_{1})_{*}$ that maps an initial distribution $\mu_{0}$ to the probability measure $\mu_{n}=(T_{n}\cdots T_{2}T_{1})_{*}\mu_{0}$ modeling the probability distribution of the $n^{th}$ forward iterate $\theta_{n}=T_{n}\cdots T_{2}T_{1}(\theta_{0})$ when the initial point $\theta_{0}$ is randomly sampled from $\mu_{0}$. Observe that the forward iterates $\theta_{n}$ form a sequence of random variables, which is a Markov chain (since for the forward iterates we have that $\theta_{n}=T_{n}(\theta_{n-1})$). Many previous works (see related work in Section˜1.1) have shown under different assumptions that this sequence of forward iterates do not converge to a point when using fixed learning rates, but rather their probability distributions $\mu_{n}$ converge to a limiting probability distribution, called the stationary distribution of the Markov chain. In other words, starting at initial point $\theta_{0}$ the distribution $\mu_{n}=(T_{n})_{*}\cdots(T_{1})_{*}\delta_{\theta_{0}}$ of the forward iterate $\theta_{n}$ converges to a stationary distribution $\mu_{\theta^{*}}$. Recall that $\delta_{\theta_{0}}$ is the Dirac measure concentrated at $\theta_{0}$ (i.e., $\delta_{\theta_{0}}(A)$ is $1$ if $A$ contains $\theta_{0}$ and 0 otherwise).
On the contrary, as we have seen in Theorem˜2.2, the backward iterates converge to actual points in regions where the maps $T_{i}$’s become contractions (which we will call contractive regions). The following theorem (whose proof is in Appendix˜E) shows the relation between the point-wise convergence of the backward iterates and the distributional convergence of the forward iterates:
@@
Consider a sequence $\{T_{i}\}$, $i=1,2,\dots$ of independent and identically distributed random operators. Suppose that for $\theta_{0}\in\Omega=\mathbb{R}^{d}$ the backward iterates converge to a random point (randomness is due to the sampling of the random operators $T_{i}$’s):
@@ 3.1 Example: Quadratic loss
Consider the quadratic loss function $L(\theta)=\theta^{2}/2$ with $\theta\in\mathbb{R}$. Its gradient is $\theta$. We want to find the minimum of $L$ (which of course is 0 at $\theta=0$) by a stochastic gradient descent. To model the batch noise, we add an i.i.d random error $\epsilon_{i}$ to the gradient at step $i$, yielding the iteration procedure:
@@
$T_{1}\cdots T_{n}(\theta)=(1-h)^{n}\theta+h\epsilon_{1}+\cdots+h(1-h)^{n-1}\epsilon_{n}.$
@@ 3.2 Example: Learning dynamics close to minima
$L(\theta)=\frac{1}{N}\sum_{i=1}^{N}L_{i}(\theta),$
@@
since at a minimum $g(\theta^{*})=0$. To simplify the notation, we will denote by $g_{i}(\theta)$ the gradient of the batch loss $L_{i}(\theta)$, which is now a random vector (because the batch is random). We can model this stochastic gradient as follows: $g_{i}(x)=g(x)-\epsilon_{i}$, where we assume that the differences $\epsilon_{i}=g(x)-g_{i}(x)$ are i.i.d random vectors with zero mean. Using gradient approximation ˜5 we have that a single gradient update is given in a form very similar to that of the previous example:
$T_{i}(\theta)=\theta-h\Big(g(\theta)-\epsilon_{i}\Big)\simeq\theta^{*}+(1-hH)(\theta-\theta^{*})+h\epsilon_{i}$
@@
and since $\|1-hH\|<1$ for $h$ strictly smaller than $2/\lambda_{\mathrm{max}}$ (as in Example˜2.1), Theorem˜2.2 tells us that the backward trajectory converges. Now let us have a look at the form of the forward and backward iterates in this particular case to give us an intuition about why this is true. In a similar way than in the previous example, the $n^{th}$ forward SGD iterate is given by
$T_{n}\cdots T_{1}(\theta)=\theta^{*}+(1-hH)^{n}(\theta-\theta^{*})+h\epsilon_{n}+h(1-hH)\epsilon_{n-1}+\cdots\cdots+h(1-hH)^{n-1}\epsilon_{1}.$
While the term $(1-hH)^{n}(\theta-\theta^{*})$ converges to zero (when $h<2/\lambda_{max}$), we see that at each step $n$ a fresh random vector $h\epsilon_{n}$ is added to the iterate, making the whole sequence converge to a distribution rather than to a point. On the other hand the backward iterate is obtained by reversing the indices:
$T_{1}\cdots T_{n}(\theta)=\theta^{*}+(1-hH)^{n}(\theta-\theta^{*})+h\epsilon_{1}+h(1-hH)\epsilon_{2}+\cdots\cdots+h(1-hH)^{n-1}\epsilon_{n}.$
This expression converges as $n\rightarrow\infty$ since the terms are decreasing at an exponential rate as $n$ grows. Thus the iterates converge to a point now rather than a distribution. Looking at the actual form of the backward iterates, it is clear that they are distributed by the probability measure given by the forward iterates as prescribed by Theorem˜2.6.
@@ 4 Experiments
Note on plotting multiple seeds: We are interested in the variability per realization of the backward and forward trajectories. The backward trajectories are more stable individually along their own path, but these paths can be very different from seed to seed (because of the convergence toward different points). Therefore the phenomenon is much clearer when the seeds are plotted individually rather than averaged, which creates artificially more variability in the backward trajectories as there really is on each individual realization. This is why we are reporting the various seed plots individually and not as a single averaged plot with error bars. Note that the increased stability and point convergence is visible in each of the seeds, which are added to the Appendix˜B.
Figure 2: Backward SGD exhibits decreased variance and increased stability compared to forward SGD for a ResNet-18 model trained on CIFAR-10. The additional seeds are in Section˜B.1.
@@
We trained a ResNet-18 with stochastic gradient descent and no regularization on the CIFAR-10 dataset Krizhevsky (2009). We used a learning rate of 0.025 and a batch-size of 8. In Figure˜2, we recorded the learning curves at each gradient update for both the forward and backward iterations for 2000 steps. The additional seeds are in Section˜B.1. We observe that the training loss for the backward trajectory is more stable, converges faster, and has less variability than the forward trajectory, and similarly for all the other learning curves. In Appendix˜A, we repeat the experiment with different datasets (synthetic dataset, FashionMNIST, and CIFAR-100) as well as different architectures (ResNet-50, VGG, and MLP) and using different base optimizers (AdamW). Each time we observe the same phenomenon for all 5 seeds.
@@
We trained a MLP with 5 layers of 500 neurons each with stochastic gradient descent with no regularization to learn FashionMNIST Xiao et al. (2017). The batch-size was set to 8 while the learning rate was 0.001. In Figure˜3, we report the test accuracy at every single step for 2000 steps. At step 1000, we reset the point from which we perform backward SGD to that of the parameter at this step. As a result we see that the backward trajectory from that reset point seems to converge again, but to a different point (with different test accuracy), while forward SGD seems to oscillate between these two points. This is in line with the theoretical prediction that backward trajectories SGD converge toward points distributed according to the distribution induced by forward trajectories (see Theorem˜2.6). The additional seeds are in Section˜B.2.
Figure 3: Backward SGD converges toward a different minimum after resetting the initialization point at step 1000 (“intermittent backward”) while forward SGD oscillates between them for MLP trained on FashionMNIST. Top: On the first seed, backward changes from a higher test-performance trajectory to a lower test-performance trajectory at the reset step 1000. Bottom: On the second seed, backward changes this time from a trajectory converging to a lower test-performance point to a trajectory converging to a higher test-performance point. The other seeds can be found in Section˜B.2 including all the learning curves.
@@ 5 Conclusion and future work
In this paper, we analyzed the impact of two different composition orders (to produce each iterate) on the convergence and stability of stochastic learning trajectories. Although the very idea of backward trajectories may seem strange and counterintuitive at first, we show that this approach of backward optimization has clear advantages. Namely, we show theoretically that the backward trajectories converge toward actual points while the forward trajectories converge toward probability distributions in regions where the optimization maps become contractions, leading to improved stability and convergence. We observed this phenomenon experimentally on a number of datasets (synthetic, FashionMNIST, CIFAR-10, and CIFAR-100) and neural network architectures (MLP, VGG-19, ResNet-18, and ResNet-50), making it relevant to deep learning. This phenomenon points toward the fact that the ordering in which the data examples are consumed to produce each parameter iterate impacts the properties of both the learning trajectories and their convergence points. In particular the last examples used to produce a parameter iterate seem to have great importance. Indeed, the backward order is the only order in which the sequence of the last examples used to produce a parameter iterate remains consistent at each training step. As we saw, this yields convergence toward a point rather than oscillations between solutions of varied performances.
@@
One possible strategy to mitigate the intensive computational requirements of the full backward order is to apply it only partially on a fixed window of iterations from a point that is reset periodically. This way the extra computation time is capped by the window size. We see that this approach leads to convergence and increased stability in Figure˜3; however the backward trajectory may change and converge to a different point at each reset. We can imagine strategies to initiate this reset only when the performance of the new backward trajectory increases, leading to stable convergence to higher test performance points. Another possibility is to start the backward order only later in training to force convergence toward a single solution after the forward trajectory has done enough exploration of the space as demonstrated in Figure˜11 in Appendix˜A. We imagine that this approach may possibly lead toward more reliable early-stopping criteria. Similarly, application of the backward order at the very beginning of training may give an accurate idea of whether a specific hyperparameter setting is promising or needs to be aborted by creating stabler trajectories whose learning curves are easier to interpret early on.
An orthogonal approach to deal with the computational challenges of backward optimization is to approximate the backward trajectory. In Appendix˜C, we show how to compute a backward iterate from a cheaper forward iterate to which we add a correction that can be computed completely independently. We compute these corrections up to order 2 in Theorem˜C.2 Theorem˜C.2, namely:
@@
where $\nabla L_{i},\nabla L_{j}](\theta)=H_{i}(\theta)\nabla L_{j}(\theta)-H_{j}(\theta)\nabla L_{i}(\theta)$ is the Lie bracket between the vector fields $\nabla L_{i}$ and $\nabla L_{j}$. Even though the approximation we provide is not immediately useful because the higher order terms seem to matter, it probably will be useful if higher order corrections are included. As a proof of the usefulness of these type of corrections added to the forward iterates to emulate alternate orderings, we show in Appendix˜D that the corrections we computed at second order are already enough to approximate an average between all the possible iteration orders on a batch window. Adding these corrections to the forward iterates produces a beneficial regularizer that mimics small batch training regularization. In fact, a training step can be seen as a parameter average of mixture of models where each model is trained not from a different seed but from a different batch order. We hope that this paper will bring awareness and foster more research toward understanding and exploiting the role of iteration order in the production of models with more consistent properties along more stable learning trajectories.
Broader Impact Statement
This work focused on the theoretical aspect of learning algorithms, especially stability. We do not foresee any negative impact from this theoretical work.
Acknowledgments
The third-listed author was supported by the Swiss NSF Grants 200020-200400 and 200021-212864, and by the Swedish Research Council Grant 104651320. We are grateful to Massimo Horvath for several insightful comments on an earlier version of the manuscript. We would like to thank Michael Wunder, Peter Bartlett, George Dahl, Atish Agarwala, and Scott Yak for helpful discussion and feedback.